OpenStudy (anonymous):

A student is solving the equation 2^x – 1 = 16^x + 2. Step 1: 2^x – 1 = 16^x + 2 Step 2: 2^x – 1 = (2^4)^x + 2 Step 3: 2^x – 1 = 2^4x + 8 Which is the next step?

OpenStudy (anonymous):

2(x 1) = 4(x + 2) x - 1 = 4x + 8 x - 1 = x + 2 x - 1 = (4x + 8)

OpenStudy (barrycarter):

In your original question, do you mean "2^x-1" or "2^(x-1)"?

OpenStudy (anonymous):

@barrycarter 2^(x-1)